Qd = 960 – 30P ; Qs = 60 + 60P *


1
Expert's answer
2022-04-04T09:14:54-0400

Q"_d" = 960 – 30P

Q"_s"= 60 + 60P

At Equilibrium, "Q_s= Q_d"

960 – 30P=60 + 60P

90P= 900

P= 10

Q= 960-30(10)

= 960-300

= 660